OpenStudy (anonymous):

(2/x - 3/y) / (2/x + 3/y) simplified

OpenStudy (campbell_st):

start by getting common denominators \[\frac{2}{x} - \frac{3}{y} = \frac{2y -3x}{xy}\] same for the denominator... then flip and multiply... look for common factors that can be removed.

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

is the denominator poitive though?

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

\[\frac{\frac{2}{x}-\frac{3}{y}}{\frac{2}{x}+\frac{3}{y}}=\frac{\frac{-3 x+2 y}{x y}}{\frac{3 x+2 y}{x y}}=\frac{-3 x+2 y}{3 x+2 y}=\frac{2 y-3 x}{3 x+2 y} \]

OpenStudy (anonymous):

Some of the terms may not be in classroom order.

OpenStudy (anonymous):

thats fine i got it thanks

OpenStudy (anonymous):

And thank you for the medal.